The test statistic of zequals=2.75 is obtained when testing the claim that pnot equals≠0.624

a. Identify the hypothesis test as being​ two-tailed, left-tailed, or​ right-tailed.

b. Find the​ P-value.

c. Using a significance level of alphaαequals=0.100.10​, should we reject Upper H 0H0 or should we fail to reject Upper H 0H0​?
Answer #1

solution

this is two-tailed test

P(z >2.75 ) = 1 - P(z < 2.75) = 1-0.997=0.0030

P-value = 0.0030

= 0.10

P-value <

Reject the null hypothesis .